# Mechanical hysterons with tunable interactions of general sign

### v1

@@ Abstract
Hysterons are the basic units of hysteresis that underlie many of the complex behaviors of disordered matter. Recent work has sought to develop designs for mechanical hysterons, both to better understand their interactions and to create materials that respond to their mechanical environment in novel ways. Elastic structures including slender beams, creased sheets, and shells offer the requisite bistability for artificial hysterons, but producing and controlling interactions between such structures has proven challenging. Here we report a mechanical hysteron composed of rigid bars and linear springs, which has controllable properties and tunable interactions of general sign. We derive an approximate mapping from the system parameters to the hysteron properties, and we show how collective behaviors of multiple hysterons can be targeted by adjusting geometric parameters on the fly. Our results provide a basic step towards hysteron-based materials that can sense, compute, and respond to their mechanical environment.
When an amorphous solid, a disordered magnet, or a crumpled sheet is subjected to an oscillating global drive, it navigates a complex pathway through a multitude of locally-stable states [1]. A powerful approach for understanding these systems has been to focus on the localized hysteretic rearrangements—hysterons—that arise within them [2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. In a recent model [12, 13, 14, 15], each hysteron $i$ possesses a pair of thresholds $\gamma_{i}^{+}$ and $\gamma_{i}^{-}$ at which it switches between its two possible states, $s_{i}=\pm 1$. These thresholds depart from their “bare” values, $\gamma_{i,0}^{\pm}$ due to interactions with other hysterons:
$\gamma_{i}^{\pm}(S)=\gamma_{i,0}^{\pm}-\frac{1}{2}\sum_{j\neq i}J_{ij}s_{j}\ ,$ (1)
@@
In addition to offering a foothold for understanding disordered media, Eq. 1 suggests a design principle for a class of synthetic matter. If the thresholds and interactions can be designed, materials could be created with tailored responses to their mechanical environment. As the states $s_{i}$ are readily interpreted as bits, such materials are endowed with digital memory and the ability to perform computations [16]. Several mechanical metamaterials have been developed to this end, including origami bellows [17], corrugated sheets [18], buckled beams [19, 20], and biholar sheets [21, 22]. Yet, even in this realm where one has complete control over the material architecture, no single system has been able to realize the full generality of Eq. 1.
@@
Here we develop a design for interacting mechanical hysterons, which we implement using rigid bars, low-friction bearings, and linear springs (Fig. 1). Our design enables: (i) tunable switching thresholds for each hysteron, (ii) tunable interactions between the hysterons that can be cooperative or frustrated, as well as reciprocal or non-reciprocal, and (iii) a tunable coupling to a global mechanical drive. We rationalize these results using an exact kinematic model and a simplified model that yields a mapping from the system parameters to the general model, Eq. 1. We then demonstrate the flexibility of the design by experimentally realizing all possible transition graphs for two interacting hysterons with $J_{12}J_{21}\geq 0$, and then demonstrating a “latching” behavior [23] that requires non-reciprocal interactions. Finally, we use larger collections of hysterons to accomplish more complex tasks: distinguishing between even and odd driving cycles [12] and counting the total number of driving cycles. Our results provide a basic step towards the design of materials that navigate through targeted sequences of internal states, in a way that is responsive to their mechanical environment, both present and past.
Hysteron design.— Figure 1A illustrates the basic bistable unit in our system: a rigid bar that is free to rotate about a central pivot. The bar is corralled between two posts that restrict its angle to an interval, $[\theta^{-},\theta^{+}]$, with $\theta=0$ corresponding to vertical. We achieve precise, low-friction rotation using a radial steel ball bearing with the bearing grease removed. A spring of stiffness $k_{1}$ and rest length $x_{k_{1},0}$ attaches to the bar a distance $L$ from the pivot; this spring connects to a long steel rod that translates freely along its axis via two linear bearings. The $x$-position of this rod, $\gamma$, will serve as the global mechanical drive. All springs used in the experiments behave linearly, which we verify by measuring their force-versus-displacement curves when pulling up on a known mass resting on an electronic balance (Fig. 1B).
@@
Figure 2: Transition thresholds for an interacting hysteron pair. A. Using the configuration shown in the left panel of Fig. 1D, we measure the eight transition thresholds as the mounting position of the coupling springs, $\ell$, is varied. The experiments are in good agreement with the theory. Geometric parameters: $y=9.4$ cm, $L=6.3$ cm, $w=15.2$ cm, $x_{1}=0$ cm, $x_{2}=-1.9$ cm, $[\theta_{1}^{-},\theta_{1}^{+}]=[-36.6^{\circ},36.7^{\circ}]$, and $[\theta_{2}^{-},\theta_{2}^{+}]=[-9.75^{\circ},11.2^{\circ}]$. Driving springs: $k_{1}=k_{2}=0.19$ N/cm, rest lengths $3.9$ and $4.2$ cm. Coupling springs: $k_{1,2}=0.07$ N/cm, rest lengths $2.4$ and $2.6$ cm. B. Transition graphs showing behavior for $\ell<4.18$ cm and $4.904.18<\ell<$ 4.90 cm. C. Splitting between pairs of switching thresholds, $\gamma_{i}^{\pm}(+)-\gamma_{i}^{\pm}(-)$, which indicates the strength of interactions between hysterons. Triangles: Experiments. Solid lines: Kinematic model. Dashed lines: Eq. 5 for rotors confined to small angles.
@@
The $\gamma$ that solve this equation are the switching thresholds, $\gamma_{i}^{\pm}(S)$, wherein hysteron $i$ switches out of the state $s_{i}=\mp 1$  [^2] We do not presently have a precise condition on the parameters that would prohibit stable equilibria within the interval $(\theta_{i}^{-},\theta_{i}^{+})$..
@@
Mapping to general model at small rotor angles.— A closer study of Fig. 2A reveals some interesting facts about the system. First, the interactions may be non-reciprocal, meaning that the effect of hysteron 1 on hysteron 2 is not the same as the effect of hysteron 2 on hysteron 1. This can be seen by comparing the splitting of the blue curves for hysteron 2 as compared with the black curves for hysteron 1. Second, the mean of these pairs of thresholds, $[\gamma_{i}^{\pm}(+)+\gamma_{i}^{\pm}(-)]/2$, may drift as a result of the interaction. In principle, one could understand these features of the system by examining the kinematic model, but in practice its nonlinear form conceals an intuitive understanding of these behaviors.
@@
$J_{ij}$ $=$ $4\frac{k_{ij}}{k_{i}}\frac{\ell^{2}}{L}g_{ij}\Delta\theta_{j}\ .$ (5)
Note that the sum in Eq. 4 represents the drift of the bare threshold, $\gamma_{i,0}^{\pm}$, due to the presence of interactions. This term helps preserve the up-down symmetry of the interactions in Eq. 5, where $J_{ij}=J_{ij}^{+}=J_{ij}^{-}$.
Figure 4: Suite of behaviors for an interacting hysteron pair. A. Transition graphs obtained in the experiment, as shown in Supplementary Movies 1-4. Graphs $i$ and $ii$ are “Preisach” graphs that do not require interactions [14], whereas graphs $iii$-$vi$ require cooperative interactions. To show how one can move between different transition graphs, we demonstrate them all using cooperative interactions, without changing the driving or coupling springs (“Configuration 1” in C). Graphs $vii$-$ix$ require frustrated interactions; we obtain them with a second set of driving and coupling springs (“Configuration 2” in C). B. Switching thresholds calculated from the model, using the experimental parameters. C. Experimental parameters. We fix $x_{1}=0$ cm, $y=14.5$ cm, $L=6.3$ cm, $\ell=4.7$ cm, and $w=15.2$ cm. D. Schematic illustrating the role of interactions on hysteron 2 in graphs $i$-$ix$. Top row: hysteron 1 is (+)(+)( + ); bottom row: hysteron 1 is ()(-)( - ). Smaller cooperative interactions are required to create graphs $i$-$vi$ in B, whereas larger frustrated interactions are needed for graphs $vii$-$ix$.
The non-reciprocity of the interaction can now immediately be seen, as $J_{ij}$ depends on the angular interval $2\Delta\theta_{j}=\theta_{j}^{+}-\theta_{j}^{-}$ but not on $\theta_{i}$. The larger splitting of hysteron 2 in Fig. 2A is thus due to the larger angular interval of hysteron 1. To test this result quantitatively, Fig. 2C compares Eq. 5 with the value of $\gamma_{i}^{\pm}(+)-\gamma_{i}^{\pm}(-)$ from the full kinematic model and the experiments. The comparison shows that Eq. 5 captures this non-reciprocity quantitatively, as well as the quadratic dependence of $J_{ij}$ on $\ell$.
To test the small-$\theta$ model further, Fig. 3 compares the kinematic model with the small-$\theta$ model for a pair of hysterons with parallel springs (Fig. 3A) and crossed springs (Fig. 3B). The parameters are otherwise the same as in Fig. 2. The small-$\theta$ model captures the gross trends of all 8 switching thresholds in both configurations. In particular, the drift in $[\gamma_{i}^{\pm}(+)+\gamma_{i}^{\pm}(-)]/2$ is captured by the small-$\theta$ model, and is due to the last term in Eq. 4, which measures the difference between the mean rotor angle $\overline{\theta_{j}}$ and the current rotor angle $\theta_{i}$ [recall that the ($+$) transition is out of the (--) state and vice versa]. This drift is the most apparent in the thresholds for hysteron 1 in Fig. 3, as that rotor is bounded by larger angles. In general, the pair $\gamma_{i,0}^{+}(s_{j})$ and the pair $\gamma_{i,0}^{-}(s_{j})$ drift in opposite directions, with the sign of the drift set by $g_{ij}$.
@@
Latching.— The ability to manifest non-reciprocal frustrated interactions in this system allows one to access a behavior termed “latching”. This behavior was proposed by Lindeman et al. [23] to explain how amorphous solids form memories of asymmetric driving. Here we demonstrate it with our mechanical hysterons in Fig. 5A and Supplementary Movie 5. Starting in the --- - state at $\gamma=0$, we drive the system to $\gamma=10$ cm, which puts the hysterons in the ++-+ - state. This state remains stable as we return $\gamma$ back to 00; hysteron 1 is now “latched” in the (+)(+)( + ) state. A larger driving amplitude to $\gamma=32$ cm and back to 00 releases hysteron 1 back to ()(-)( - ) via a sequence of intermediate states: ++++,+$\rightarrow$+$\rightarrow$-+$\rightarrow$--+ - → + + → - + , → - -. This behavior necessarily violates return-point memory [26, 27, 28], which would stipulate that the same state would recur each time the driving returns to its common endpoint, $\gamma=0$.
@@
Lindeman et al. [23] showed how latching is tied to the ordering of the switching thresholds; we now show how we translate this ordering into a concrete design in our mechanical system. First, latching requires $0<\gamma_{2}^{-}(+)-\gamma_{2}^{-}(-)<\gamma_{1}^{-}(+)-\gamma_{1}^{-}(-)$. This condition signals a non-reciprocal frustrated interaction, which we create by confining hysteron 1 to a smaller angular interval than hysteron 2 and coupling them together with crossed springs. Second, latching requires $\gamma_{2}^{-}(+)<\gamma_{1}^{-}(+)$ to reach the intermediate +-+- + state. To achieve this, we adjust $x_{2}$ until $\gamma_{2}^{-}(+)<\gamma_{1}^{-}(+)$, which moves all thresholds for hysteron 2 in concert, without altering their spacing. As a result, our design has $\gamma_{1}^{-}(-)<\gamma_{2}^{-}(-)$, another requirement for latching. Third, latching requires $\gamma_{1}^{-}(-)<0<\gamma_{2}^{-}(-)$ so that the state +-+- + is released as $\gamma\rightarrow 0$ but the state ++-+ - is not. We fulfill this requirement by adjusting $x_{1}$. Finally, we need $\gamma_{1}^{+}(-)<\gamma_{2}^{+}(-)$; this turns out to be satisfied in the design through the larger angular interval of hysteron 2.
@@
Figure 5B shows the switching thresholds as measured in our experimental realization of latching, using the parameters in the caption. Plugging these parameters into the kinematic model, we obtain the second set of bars, which are in good agreement with the experiment and follow the same ordering.
@@
Our mechanical system allows one to latch and release the hysterons in another way, shown in Supplementary Movie 6. Holding the driving rod fixed at $\gamma=0$, the left rotor can be flipped between ()(-)( - ) the (+)(+)( + ) by hand, and it is stable in either state when the right rotor is ()(-)( - ). Flipping the right rotor to (+)(+)( + ) immediately releases the left rotor to ()(-)( - ), returning the system to --- - upon letting go.
Figure 6: Counting modulo 2 with four mechanical hysterons. A. Thresholds and interactions that yield a limit cycle with twice the period of the driving, for unit-amplitude cyclic drive [12]. B. Corresponding pathway, where we have colored the two states at the minimum of the driving. For simplicity we do not show all the transitions out of all the states, although they can be determined from the information in A. C. Realizing this behavior with mechanical hysterons. The rotors from left to right correspond to the hysterons from top to bottom in A. The springs connecting rotors 1 and 3 have $k=0.07$ N/cm, and their original $x_{0}=2.5$ cm are extended by steel wire. All other springs have $k=0.19$ N/cm and $x_{0}=4.1$ cm. D. States and transitions in the experiment, which achieves the targeted response. Horizontal segments denote avalanches.
@@
Counting modulo 2.— More sophisticated functions can be realized with more hysterons. Here we target a multiperiodic response that was obtained for simulated hysterons [12, 13, 15]. We start with a design from one of those studies [12], where four hysterons are coupled as in Fig. 6A. When cyclic driving of unit amplitude is applied [29], these parameters produce the steady state drawn in Fig. 6B, which we find using the open-source hysteron software package [29]. The sequence of states takes two driving cycles to traverse. These interacting hysterons thus perform a basic computation: discriminating between even and odd numbers of driving cycles, i.e., counting modulo 2.
@@
Investigating the pathway in Fig. 6B further, we can see that it features two avalanches (+-+- ++-+- + $\rightarrow$ +++++ ++++++ + and +++++ $+$-+ - $\rightarrow$ --- -++-+ -), as well as a pair of “scrambled” transitions: +++++ ++++++ + $\rightarrow$ +++++ $+$-+ - and --- -+++++ + $\rightarrow$ --- -+-+- +, meaning that they are not consistent with a state-independent ordering of the switching fields [18, 14]. To see this, note that these transitions imply that $\gamma_{3}^{-}>\gamma_{4}^{-}$ when $s_{1}=s_{2}=-1$, and that $\gamma_{3}^{-}<\gamma_{4}^{-}$ when $s_{1}=s_{2}=1$. The presence of either scrambling or avalanches is known to be necessary, but not sufficient, for a multiperiodic response [14].
Figure 7: Counting driving cycles with a chain of mechanical hysterons. Time series showing a domain wall ratcheting down an antiferromagnetic chain, irreversibly flipping one hysteron each half cycle. Here a cycle of driving follows $\gamma:\gamma_{0}\rightarrow-\gamma_{0}\rightarrow\gamma_{0}$. The growing +-+- ++-+- +... phase is highlighted in blue. For ease of construction, the radial bearings are replaced by optical posts in post holders lubricated with silicone oil. Coupling springs: $k=0.17$ N/cm, $x_{0}=4.0$ cm. Driving springs: $k=0.19$ N/cm, $x_{0}=4.1$ cm.
@@
Counting to $n$.— As a final demonstration of computing with mechanical hysterons, Fig. 7 and Supplementary Movie 8 show an accumulator, which counts the number of driving cycles applied to it. A mechanical metamaterial was recently designed with this function [30], using a train of carefully designed elastic beams that make and break contact under cyclic compression. Our design is based on an antiferromagnetic chain, here a series of hysterons with uniform frustrated interactions. We stagger the points where the springs attach to the driving rod, which alternatingly raises and lowers the values of $\gamma^{+}$ and $\gamma^{-}$ for each hysteron. In particular, we set $x_{2m-1}=-L\sin(\theta_{0})$, $x_{2m}=L\sin(\theta_{0})$, and $x_{1}=-3L\sin(\theta_{0})$, where $\theta_{0}=\theta_{i}^{+}=-\theta_{i}^{-}$, although other values can also work. This design breaks the symmetry of the ground state so that ++-+ -++-+ -++-+ - has a higher energy than +-+- ++-+- ++-+- +.

### v2

@@ Abstract
Hysterons are elementary units of hysteresis that underlie many complex behaviors of non-equilibrium matter. Because models of interacting hysterons can describe disordered matter, this suggests that artificial systems could respond to mechanical inputs in precise and targeted ways. Specifying the properties of hysterons and their interactions could thus be a general method for realizing arbitrary non-equilibrium behaviors. Elastic structures including slender beams, creased sheets, and shells are clear candidates for artificial hysterons, but complete control of their interactions has seemed impractical or impossible. Here we report a mechanical hysteron composed of rigid bars and linear springs, which has controllable properties and tunable interactions of general sign that can be reciprocal or non-reciprocal. We derive a mapping from the system parameters to the hysteron properties, and we show how collective behaviors of multiple hysterons can be targeted by adjusting geometric parameters on the fly. By transforming an abstract hysteron model into a physical design platform, our work demonstrates a route toward designed materials that can sense, compute, and respond to their mechanical environment.
When an amorphous solid, a disordered magnet, or a crumpled sheet is subjected to an oscillating global drive, it navigates a complex pathway through a multitude of locally-stable states [1]. A powerful approach for understanding these systems has been to focus on the localized hysteretic rearrangements—hysterons—that arise within them [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. In such models, each hysteron $i$ possesses a pair of thresholds $\gamma_{i}^{+}$ and $\gamma_{i}^{-}$ at which it switches between its two possible states, $s_{i}=\pm 1$. More general models incorporate interactions between hysterons [13, 14, 15, 16], so that the switching thresholds $\gamma_{i}^{\pm}$ depart from their “bare” values $\gamma_{i,0}^{\pm}$ as follows:
$\gamma_{i}^{\pm}(S)=\gamma_{i,0}^{\pm}-\sum_{j\neq i}J_{ij}s_{j}\ ,$ (1)
@@
In addition to offering a foothold for understanding disordered media, Eq. 1 suggests a design principle for a class of synthetic matter. If the thresholds and interactions can be designed, materials could be created with tailored responses to different mechanical environments. As the states $s_{i}$ are readily interpreted as bits, such materials are endowed with digital memory and the ability to perform computations [17]. Several mechanical metamaterials have been developed to this end, including origami bellows [18], corrugated sheets [19], buckled beams [20, 21], and biholar sheets [22, 23]. Yet, even in this realm where one has complete control over the material architecture, no system has been able to realize the full generality of Eq. 1.
@@
Here we show that it is possible to embody this abstract hysteron model in a physical system built from simple components: linear springs, rigid rods, and low-friction bearings (Fig. 1). We design interacting mechanical hysterons that allow: (i) tunable switching thresholds for each hysteron, (ii) tunable interactions between the hysterons that can be cooperative or frustrated, as well as reciprocal or non-reciprocal, and (iii) a tunable coupling to a global mechanical drive. We rationalize these results using an exact kinematic model, with a limiting behavior that maps the experimental parameters to the thresholds and interactions of Eq. 1. We then demonstrate the flexibility of the design by realizing a broad set of transition graphs for two interacting hysterons, as well as a “latching” behavior [24] that requires non-reciprocal interactions. Finally, we use larger collections of hysterons to accomplish more complex tasks: distinguishing between even and odd driving cycles [13] and counting the total number of driving cycles.
Our findings provide a general basis for mechanisms that navigate through precise sequences of states, e.g., in robotics [25]. More broadly, our results show that the size of the design space for hysteron-based materials, and the amount of control that can be wielded within it, is perhaps larger than previously thought. Our work thus identifies a vast potential for designed materials that inherently engage with memory [26], as their current state is a by-product of their driving history along with their overall design, which our platform now brings under the experimenter’s control.
Hysteron design.— Figure 1A illustrates the basic bistable unit in our system: a rigid bar that is free to rotate about a central pivot. The bar is corralled between two posts that restrict its angle to an interval, $[\theta^{-},\theta^{+}]$, with $\theta=0$ corresponding to vertical. We achieve precise, low-friction rotation using a radial steel ball bearing with the bearing grease removed. A spring of stiffness $k_{1}$ and rest length $x_{k_{1},0}$ attaches to the bar a distance $L$ from the pivot; the other end of this spring attaches to a long steel shaft at a location $x_{1}$ along its length. The shaft translates freely along its axis via two linear bearings. The $x$-position of this shaft, $\gamma$, will serve as the global mechanical drive. All springs used in the experiments behave linearly, which we verify by measuring their force-versus-displacement curves when pulling up on a known mass resting on an electronic balance (Fig. 1B).
@@
Figure 2: Transition thresholds for an interacting hysteron pair. A. Using the configuration shown in the left panel of Fig. 1D, we measure the eight transition thresholds as the mounting position of the coupling springs, $\ell$, is varied. The experiments are in good agreement with the theory. Geometric parameters: $y=9.4$ cm, $L=6.3$ cm, $w=15.2$ cm, $x_{1}=0$ cm, $x_{2}=-1.9$ cm, $[\theta_{1}^{-},\theta_{1}^{+}]=[-36.6^{\circ},36.7^{\circ}]$, and $[\theta_{2}^{-},\theta_{2}^{+}]=[-9.75^{\circ},11.2^{\circ}]$. Driving springs: $k_{1}=k_{2}=0.19$ N/cm, rest lengths $3.9$ and $4.2$ cm. Coupling springs: $k_{1,2}=0.07$ N/cm, rest lengths $2.4$ and $2.6$ cm. B. Transition graphs showing behavior for $\ell<4.18$ cm and $4.904.18<\ell<$ 4.90 cm. C. Splitting between pairs of switching thresholds, $\gamma_{i}^{\pm}(+)-\gamma_{i}^{\pm}(-)$, which measures the strength of interactions between hysterons. Triangles: Experiments. Solid lines: Kinematic model. Dashed lines: Eq. 5 for rotors confined to small angles.
@@
The $\gamma$ that solve this equation are the switching thresholds, $\gamma_{i}^{\pm}(S)$, wherein hysteron $i$ switches out of the state $s_{i}=\mp 1$  [^2] We do not presently have a precise condition on the parameters that would prohibit additional stable equilibria within the interval $(\theta_{i}^{-},\theta_{i}^{+})$..
@@
Mapping to general model at small rotor angles.— A closer study of Fig. 2A reveals some interesting facts about the system. First, the interactions may be non-reciprocal, meaning that the effect of hysteron 1 on hysteron 2 is not the same as the effect of hysteron 2 on hysteron 1. This can be seen by comparing the splitting of the blue curves for hysteron 2 as compared with the black curves for hysteron 1. Second, the mean of these pairs of thresholds, $[\gamma_{i}^{\pm}(+)+\gamma_{i}^{\pm}(-)]/2$, may shift as a result of the interaction. In principle, one could understand these features of the system by examining the kinematic model, but in practice its nonlinear form conceals an intuitive understanding of these behaviors.
@@
$J_{ij}$ $=$ $2\frac{k_{ij}}{k_{i}}\frac{\ell^{2}}{L}g_{ij}\Delta\theta_{j}\ .$ (5)
Note that the sum in Eq. 4 represents the shift of the bare threshold, $\gamma_{i,0}^{\pm}$, due to the presence of interactions, and although it closely resembles the interaction strength $J_{ij}$ in Eq. 5, it is distinct from it. We also note that in this limit, the interactions evidently perturb the $\gamma_{i,0}^{+}$ and the $\gamma_{i,0}^{-}$ by the same amount, i.e., $J_{ij}^{+}=J_{ij}^{-}=J_{ij}$ as in Eq. 1.
The non-reciprocity of the interaction can now immediately be seen, as $J_{ij}$ depends on the angular interval $2\Delta\theta_{j}=\theta_{j}^{+}-\theta_{j}^{-}$ but not on $\theta_{i}$. The larger splitting of hysteron 2 in Fig. 2A is thus due to the larger angular interval of hysteron 1. To test this result quantitatively, Fig. 2C compares Eq. 5 with the value of $\gamma_{i}^{\pm}(+)-\gamma_{i}^{\pm}(-)$ from the full kinematic model and the experiments. The comparison shows that Eq. 5 captures the observed non-reciprocity, and that its dependence on $\ell$ is also in good agreement.
To test the small-$\theta$ model further, Fig. 3 compares the kinematic model with the small-$\theta$ model for a pair of hysterons with parallel springs (Fig. 3A) and crossed springs (Fig. 3B). The parameters are otherwise the same as in Fig. 2. The small-$\theta$ model captures the gross trends of all 8 switching thresholds in both configurations. In particular, the shift in $[\gamma_{i}^{\pm}(+)+\gamma_{i}^{\pm}(-)]/2$ is captured by the small-$\theta$ model, and is due to the last term in Eq. 4, which measures the difference between the mean rotor angle $\overline{\theta_{j}}$ and the current rotor angle $\theta_{i}$ [recall that the ($+$) transition is out of the (--) state and vice versa]. This shift is the most apparent in the thresholds for hysteron 1 in Fig. 3, as that rotor is bounded by larger angles. In general, the pair $\gamma_{i,0}^{+}(\pm)$ and the pair $\gamma_{i,0}^{-}(\pm)$ shift in opposite directions, with the sign of the shift set by $g_{ij}$.
@@
Figure 4: Suite of behaviors for an interacting hysteron pair. A. Transition graphs obtained in the experiment, showing all states and transitions accessible from the initial state, ---- (Supplementary Movies 1-4). Graphs $i$ and $ii$ are “Preisach” graphs that do not require interactions [15], whereas graphs $iii$-$vi$ require cooperative interactions. To show how one can move between different transition graphs, we demonstrate them all using cooperative interactions, without changing the driving or coupling springs (“Configuration 1” in C). Graphs $vii$-$ix$ require frustrated interactions; we obtain them with a second set of driving and coupling springs (“Configuration 2” in C). B. Switching thresholds calculated from the model, using the experimental parameters. C. Experimental parameters. We fix $x_{1}=0$ cm, $y=14.5$ cm, $L=6.3$ cm, $\ell=4.7$ cm, and $w=15.2$ cm. D. Schematic illustrating the role of interactions on hysteron 2 in graphs $i$-$ix$. Top row: hysteron 1 is (+)(+)( + ); bottom row: hysteron 1 is ()(-)( - ). Smaller cooperative interactions are required to create graphs $i$-$vi$ in B, whereas larger frustrated interactions are needed for graphs $vii$-$ix$.
@@
Latching.— The ability to manifest non-reciprocal frustrated interactions in this system allows one to access a behavior termed “latching”. This behavior was proposed by Lindeman et al. [24] to explain how amorphous solids form multiple memories of asymmetric driving. Here we demonstrate it with our mechanical hysterons in Fig. 5A and Supplementary Movie 5. Starting in the --- - state at $\gamma=0$, we drive the system to $\gamma=10$ cm, which puts the hysterons in the ++-+ - state. This state remains stable as we return $\gamma$ back to 00; hysteron 1 is now “latched” in the (+)(+)( + ) state. A larger driving amplitude to $\gamma=32$ cm and back to 00 releases hysteron 1 back to ()(-)( - ) via a sequence of intermediate states: ++++,+$\rightarrow$+$\rightarrow$-+$\rightarrow$--+ - → + + → - + , → - -. This behavior necessarily violates return-point memory [29, 30, 26], which would stipulate that the same state would recur each time the driving returns to its common endpoint, $\gamma=0$.
Lindeman et al. [24] showed how latching is tied to the ordering of the switching thresholds. They found that latching requires:
$\gamma_{1}^{-}(-)<0<\gamma_{2}^{-}(-)<\gamma_{2}^{-}(+)<\gamma_{1}^{-}(+)\ ,$ (6a)
$\gamma_{1}^{+}(-)<\gamma_{2}^{+}(-)\ .$ (6b)
@@
We now show how we translate this ordering into a concrete design in our mechanical system. We start by noting that Eq. 6a implies: $0<\gamma_{2}^{-}(+)-\gamma_{2}^{-}(-)<\gamma_{1}^{-}(+)-\gamma_{1}^{-}(-)$, which we rewrite as: $0<-J_{21}<-J_{12}$. We obtain this non-reciprocal frustrated interaction by confining hysteron 1 to a smaller angular interval than hysteron 2 and coupling them together with crossed springs following Eq. 5. Next, we note that adjusting $x_{2}$ moves all thresholds for hysteron 2 in concert without altering their spacing. We thus move $x_{2}$ until $\gamma_{2}^{-}(+)<\gamma_{1}^{-}(+)$ and $\gamma_{1}^{-}(-)<\gamma_{2}^{-}(-)$. Then, we fulfill $\gamma_{1}^{-}(-)<0<\gamma_{2}^{-}(-)$ by moving $x_{1}$ and $x_{2}$ together by equal amounts until those thresholds straddle $\gamma=0$. Finally, $\gamma_{1}^{+}(-)<\gamma_{2}^{+}(-)$ turns out to be satisfied already through the larger angular interval of hysteron 2.
@@
Figure 5B shows the switching thresholds as measured in our experimental realization of latching, using the parameters in the caption. The ordering of thresholds indeed follow Eq. 6. Plugging the experimental parameters into the kinematic model, we obtain the second set of bars, which are in good agreement with the experiment and follow the same ordering.
@@
Our mechanical system allows one to latch and release the hysterons in another way, shown in Supplementary Movie 6. Holding the driving rod fixed at $\gamma=0$, the left rotor can be flipped between ()(-)( - ) the (+)(+)( + ) by hand, and it is stable in either state when the right rotor is ()(-)( - ). From the state ++-+ -, flipping the right rotor to (+)(+)( + ) immediately releases the left rotor to ()(-)( - ), returning the system to --- - upon letting go.
@@
Counting modulo 2.— More sophisticated functions can be realized with more hysterons. Here we target a multiperiodic response that was obtained for simulated hysterons [13, 14, 16]. We start with a design from one of those studies [13], where four hysterons are coupled as in Fig. 6A. When cyclic driving of unit amplitude is applied [31], these parameters produce the steady state drawn with black and red arrows in Fig. 6B, which we determine using the open-source hysteron software package [31]. The sequence of states takes two driving cycles to traverse. These interacting hysterons thus perform a basic computation: discriminating between even and odd numbers of driving cycles, i.e., counting modulo 2.
Figure 6: Counting modulo 2 with four mechanical hysterons. A. Thresholds and interactions that yield a limit cycle with twice the period of the driving, for unit-amplitude cyclic drive [13]. B. Corresponding pathway, where we have colored the two states at the minimum of the driving. Lighter arrows show all other transitions available to the system under arbitrary driving and are not used in the period-2 cycle. States without arrows to them are unstable at all $\gamma$. C. Realizing this pathway with mechanical hysterons. The rotors from left to right correspond to the hysterons from top to bottom in A. The springs connecting rotors 1 and 3 have $k=0.07$ N/cm, and their original $x_{0}=2.5$ cm are extended by steel wire. All other springs have $k=0.19$ N/cm and $x_{0}=4.1$ cm. D. States and transitions in the experiment, which achieves the targeted response. Horizontal segments denote avalanches.
@@
Investigating this pathway in Fig. 6B further, we can see that it features two avalanches (+-+- ++-+- + $\rightarrow$ +++++ ++++++ + and +++++ $+$-+ - $\rightarrow$ --- -++-+ -), as well as a pair of “scrambled” transitions: +++++ ++++++ + $\rightarrow$ +++++ $+$-+ - and --- -+++++ + $\rightarrow$ --- -+-+- +, meaning that they are not consistent with a state-independent ordering of the switching fields [19, 15]. To see this, note that these transitions imply that $\gamma_{3}^{-}>\gamma_{4}^{-}$ when $s_{1}=s_{2}=-1$, and that $\gamma_{3}^{-}<\gamma_{4}^{-}$ when $s_{1}=s_{2}=1$. The presence of either scrambling or avalanches is known to be necessary, but not sufficient, for a multiperiodic response [15].
@@
Counting to $n$.— As a final demonstration of computing with mechanical hysterons, Fig. 7 and Supplementary Movie 8 show an accumulator, which counts the number of driving cycles applied to it. A mechanical metamaterial was recently designed with this function [32], using a train of carefully designed elastic beams that make and break contact under cyclic compression. Our design is based on an antiferromagnetic chain, here a series of hysterons with uniform frustrated interactions. We stagger the points where the springs attach to the driving rod, which alternatingly raises and lowers the values of $\gamma^{+}$ and $\gamma^{-}$ for each hysteron. In particular, we set $x_{2m-1}=-L\sin(\theta_{0})$, $x_{2m}=L\sin(\theta_{0})$, and $x_{1}=-3L\sin(\theta_{0})$, where $\theta_{0}=\theta_{i}^{+}=-\theta_{i}^{-}$, although other values can also work. This design breaks the degeneracy of the two ground states so that ++-+ -++-+ -++-+ - has a higher energy than +-+- ++-+- ++-+- +.
@@
Figure 7: Counting driving cycles with a chain of mechanical hysterons. Time series showing a domain wall ratcheting down an antiferromagnetic chain, irreversibly flipping one hysteron each half cycle. The system is initialized in the state ++-+ -++-+ -++-+ - and a cycle of driving follows $\gamma:\gamma_{0}\rightarrow-\gamma_{0}\rightarrow\gamma_{0}$. The growing +-+- ++-+- +... phase is highlighted in blue. For ease of construction, the radial bearings are replaced by optical posts in post holders lubricated with silicone oil. Coupling springs: $k=0.17$ N/cm, $x_{0}=4.0$ cm. Driving springs: $k=0.19$ N/cm, $x_{0}=4.1$ cm.